OpenStudy (anonymous):

How is each function related to y=x? give the fuction by translating the parent function y= x-3 y=x+4.5 y=x+1.5

OpenStudy (anonymous):

RECIEVE A MEDAL

jigglypuff314 (jigglypuff314):

parent function: y = x when there is number after x, then graph move up or down by that much so for y = x - 3 so the graph ^ was moved down 3 units down because of the negative. now do you understand how to go about doing the rest? :)

OpenStudy (anonymous):

@jigglypuff314 yes! thank you but how do i answer it?

jigglypuff314 (jigglypuff314):

? are you asked to graph or something?

OpenStudy (anonymous):

@jigglypuff314 no it just says to give the function by translating the parent function

jigglypuff314 (jigglypuff314):

hmm.. I thought those were functions :/

OpenStudy (anonymous):

@jigglypuff314 im confused lol it says how are they related to y=x

jigglypuff314 (jigglypuff314):

y= x-3 is y=x moved down 3 units

OpenStudy (anonymous):

oh ok thank you @jigglypuff314

jigglypuff314 (jigglypuff314):

glad I could help :)